### v1

Abstract
We present a novel implementation of a soft sphere, discrete elements code to simulate the dynamics of self-gravitating granular materials. The code is used to study the outcome of sub-sonic collisions between self-gravitating rubble piles with masses ranging from $\sim 6\times 10^{21}$ to $\sim 6\times 10^{22}$ g. These masses are representative of asteroids and planetesimals in the $\sim 100$ km range. We simulate rubble piles composed of a range of particle sizes and analyze the collisions outcome focusing on the properties of the largest surviving fragment. We successfully test and validate the code against previous results. The results of our study show that rubble piles formed by collision of two parent rubble piles do not maintain the same particle size distribution as their parents. Rubble piles formed in low velocity collisions are characterized by a larger fraction of large particles, while the largest fragments of high-velocity collisions show a significant decrease in their mean particle size. We ascribe this effect to the fact that large particles transmit most of the forces during the collisions. In addition, we find that the mass of the largest post-collision fragment does depend on the rotation of the colliding rubble piles. This effect is especially noticeable when the pre-collision spin axes are parallel with each other and perpendicular to the relative velocity. This finding can be particularly relevant for meter to kilometer sized bodies embedded in protostellar accretion disks, where viscous stresses can efficiently align the target and projectile spin axes.
1 Introduction
While most of the baryonic matter in the Universe is in gaseous and plasma phase within stars, galaxies, and the intergalactic medium (Copi et al., 1995; Péroux & Howk, 2020), solids do play many crucial roles in the structure and evolution of the Cosmos (Draine, 2003). Solid phase materials are commonly found in the interstellar medium (ISM) as dust grains (Mathis, 1990; Weingartner & Draine, 2001; Calzetti et al., 2000) and in planetary systems as rocky planets and minor bodies, such as asteroids, comets, etc (Lissauer, 1993; Chapman et al., 1978; Walsh, 2018; Spinrad, 1987). Dust grains are known to have an effect on the thermodynamics (Spitzer, 1978) of the ISM and to be receptacles for interstellar chemistry (Duley & Williams, 1981; Hasegawa & Herbst, 1993). Collisions among grains lead to the formation of larger objects that can grow to become asteroids, planetesimals, and eventually terrestrial planets (Goldreich & Ward, 1973; Johansen et al., 2007; Blum & Wurm, 2008; Wyatt, 2008). Some of the solids in space are whole, single pieces, while others — like dust conglomerates or asteroids and planetesimals — are clusters made by loosely bound individual units or particles (Walsh, 2018). The larger clusters are held together mainly by gravity, while the small ones are dominated by contact forces.
Symbol           Description Units
$R_{i}$ Radius of the $i$-th particle cm
$m_{i}$ Mass of the $i$-th particle cm
$\vec{r}_{i}$ Position vector of the $i$-th particle cm
$r_{ij}$ Distance between the $i$-th and $j$-th particles cm
$\hat{r}_{ij}$ Unit vector that points from the $j$-th to the $i$-th particles’ positions
$\vec{v}_{i}$ Velocity vector if the $i$-th particle cm/s
$\vec{\omega}_{i}$ Angular velocity vector if the $i$-th particle rad/s
$\vec{F}_{\rm{G,ij}}$ Gravity force on particle $i$ due to particle $j$ dyne
$\vec{F}_{\rm{elastic}}$ The normal force between two particles due to the elatic reaction between the two dyne
$\vec{F}_{\rm{i,friction}}$ The kinetic friction force on particle $i$ dyne
$\vec{F}_{\rm{i,rolling\,friction}}$ The rolling friction force on particle $i$ dyne
$\vec{T}_{\rm{ij}}$ The torque on particle $i$ due to its interaction with particle $j$ dyne cm
$k$ Spring constant dyne/cm
$\mu_{k}$ Coefficient of kinetic friction
$\mu_{r}$ Coefficient of rolling friction
$\delta{t}$ time step of the code s
Table 1: List of symbols used and their meaning.
In general, based on observational evidence and previous studies (Wetherill, 1967; Dohnanyi, 1969; Chapman et al., 1978; Farinella et al., 1982; Davis et al., 1985; Michel et al., 2001; Benavidez et al., 2012; Leinhardt & Stewart, 2012; Walsh, 2018), the outcome of collisions between clusters is not always constructive. Here, by constructive we mean that the largest debris of the collision is more massive than the target cluster [^1] Here and in the following we will call the projectile the cluster with the lower mass, while we will call the target the more massive one. Since our calculations are carried out in the center of mass frame, in which the net momentum is zero, the projectile also has the largest velocity. In case the two rubble piles have the same mass, one of the two will arbitrarily be the projectile and the other the target, if such distinction is needed.. While it is expected that higher velocity collisions have a higher likelihood of being destructive, the collision outcome is not simply predictable by the energy balance of the collision. Previous work (Farinella et al., 1982; Love & Ahrens, 1996; Asphaug et al., 1998; Durda et al., 2004; Richardson et al., 2009) has indeed shown that even collisions with larger kinetic energy than binding energy (in absolute value) can be constructive, resulting in a massive, bound cluster surrounded by fast debris that carry most of the excess kinetic energy. Additional energy is dissipated in non-elastic collisions that increase the thermal energy of the particles. Leinhardt & Stewart (2012), for example, found that even collisions in which the center of mass (CoM) kinetic energy is three times larger than the gravitational potential binding energy of the clusters can be constructive. The high resilience of granular objects (or clusters) to destructive collisions can be ascribed to two main properties that differentiate them from gaseous bodies. First, clusters of hard particles obey the dynamics of granular materials, in which forces are not due to pressure gradients but to contact stresses, and form filamentary strains that propagate the impact in a disproportionate way to a few particles in the cluster, leaving the majority only marginally affected (Sánchez & Scheeres, 2011, 2012, 2014). This creates the ”fast debris” effect in which a few particles acquire high speed from the collision, in greater excess from the escape velocity. These fast particles absorb the incoming kinetic energy, allowing the bound remnant to grow more massive than the initial target cluster. In addition, since the forces generated by the collisions are propagated through the clusters via contact points, the non-elasticity of the particles that compose the cluster allow them to absorb a fraction of the impact energy at every contact point Tanga et al. (2009); Ferrari & Tanga (2020). This allows for some of the impact energy to be converted into internal energy (hotter particles).
In this paper we present a novel code to study the outcome of subsonic collisions between clusters with masses in the asteroid and planetesimal range. We first present the code (Section 2) and compare its results with the outcome of previous studies (Section 3.1). We then use the capabilities of our code to study the effect of collisions on the particle size distribution of the pre- and post-collision clusters, as well as the effect of rotation of the projectile and target clusters (Section 3). We also analyze the dynamic of the collisions to tease out which particle size is mostly involve in propagating forces during the impact. We finally discuss our results and compare with previous studies (Section 4).
2 The DECCO Code
In this section we present our soft-sphere, discrete element code DECCO (Discrete Element Cosmic COllision) that we used in the calculations presented in section 3. The code uses the soft-sphere technique to compute the binary collision between particle pairs [^2] Collisions between multiple particles are also considered as the cumulative effect of multiple binary collisions involving the same particles., and includes the effects of non-elastic collisions as well as kinetic and rotational friction between particles. At the present stage of development, DECCO only includes gravity as the attractive force and is therefore best suited for collisions between macroscopic bodies, such as rubble-pile asteroids or planetesimals. Since DECCO is based on individual particle-particle interactions, it is applicable for impact velocities that are lower than the speed of sound in the particles. Hydrodynamic codes would be needed for faster collisions (see, e.g., Michel et al. 2001; Leinhardt & Stewart 2012; Michel et al. 2020). DECCO also calculates explicitly all binary interaction, yielding increased accuracy at the cost of $O(n^{2})$ complexity.
2.1 Free particles
When there is no contact between two particles, i.e., the distance from their centers is larger than the sum of the two particle radii ($r_{ij}>Ri+R_{j}$), the only acting force is gravitational attraction. In this case the force that the $i$-th particle feels because of the presence of the $j$-th particle is:
$\vec{F}_{ij}=\vec{F_{G}}_{,ij}=-G\frac{m_{i}m_{j}}{r_{ij}^{2}}\hat{r}_{ij}$ (1)
where
$\hat{r}_{ij}=\frac{\vec{r}_{i}-\vec{r}_{j}}{r_{ij}}$ (2)
is the unit vector that points from the position of the $j$-th particle to the position of the $i$-th one.
Figure 1: Example of a pre-formed rubble-pile (left) and the same after the collapse phase, when formation is completed (right).
Figure 2: Example outcome of the post-collision sub-cluster identification algorithm. The figure shows the post-collapse configuration from an $\eta=3$ collision between two identical rubble piles (5000 particles each). The area shown is where the largest debris are found. Particles belonging to dynamically bound sub-clusters with at least 50 particle members are color coded. Dark gray particles are free and unbound. There are six sub-clusters identified in the figure. The two biggest ones (blue and green) are almost identical due to the symmetry of the collision.
2.2 Radial contact forces
The implementation of DECCO presented here includes the elastic repulsion between two touching particles as the only radial contact force. To take into account the non-elastic behaviour of real materials, DECCO utilizes two different values of Hook’s constant: a larger value $k_{\rm{in}}$ for the incoming trajectory (particles approaching each other) and a smaller value $k_{\rm{out}}$ for the outgoing trajectory. It can be easily shown that this methodology ensures a constant coefficient of restitution $\epsilon=\sqrt{k_{\rm{out}}/k_{\rm{in}}}$.
The elastic force equation therefore reads as:
$\vec{F}_{\rm{elastic}}=k\frac{R_{i}+R_{j}-||\vec{r}_{ij}||}{2}\hat{r}_{ij}$ (3)
where $||\vec{r}_{ij}||$ is the distance between particles $i$ and $j$, $R_{i}$ and $R_{j}$ are the respective radii of particles $i$ and $j$, and the factor $\frac{R_{i}+R_{j}-||\vec{r}_{ij}||}{2}$ measures the compression experienced by particle $i$ from particle $j$. When particles are colliding, this additional surface normal force is added to the gravitational force for the current timestep to simulate the collision ‘bounce’.
2.3 Restitution
As introduced above, Kinetic energy is removed from particles in collision by a coefficient of restitution directly applied through a change in the elasticity constant, $k$. This method allows for scale-free, constant restitution, which is particularly important since we aim at simulating objects made by multiple particles of different sizes.
2.4 Friction
To determine surface friction, two primary physical quantities must be derived. These are the relative surface velocity experienced by particle 1 from particle 2, and the normal force, $\vec{N}$, exerted during collision. The normal force of collision is the elastic repulsion force, $\vec{F}_{elastic}$ from Eq. 3. The relative surface velocity that particle $i$ experiences from particle $j$ depends on both particles’ angular velocity, as well as the component of their translational velocity tangent to their surfaces at the point of collision.
The surface velocity of particle $i$ at the point of contact with particle $j$ is found as:
$\vec{v}_{i,surface}=\underbrace{\vec{v}_{i}-(\vec{v}_{i}\cdot\vec{r}_{ij})\frac{\vec{r}_{ij}}{||\vec{r}_{ij}||}}_{\text{velocity component tangent to surface}}+\underbrace{\vec{\omega}_{i}\times\frac{R_{i}}{R_{i}+R_{j}}\vec{r}_{ij}}_{\text{velocity due to rotation}}$ (4)
Where $\vec{\omega_{i}}$ is angular velocity of particle $i$. The relative surface velocity is then the difference between the surface velocities of the two particles:
$\vec{v}_{i,surface,relative}=\vec{v}_{j,surface}-\vec{v}_{i,surface}$ (5)
The friction force experienced by particle $i$ is therefore
$\vec{F}_{i,friction}=-\mu_{k}||\vec{F}_{\rm{elastic}}||\hat{v}_{i,surface,relative}$ (6)
where $\mu_{k}$ is the coefficient of friction.
In addition to sliding friction, the code includes rolling friction, which is calculated as:
$\vec{F}_{i,rollingFriction}=-\mu_{r}||\vec{F}_{\rm{elastic}}||\frac{(\vec{\omega}_{i}-\vec{\omega}_{j})\times\vec{r}_{ij}}{||(\vec{\omega}_{i}-\vec{\omega}_{j})\times\vec{r}_{ij}||}$ (7)
where $\mu_{r}$ is the coefficient of rolling friction. It has to be noted that modeling rolling friction is a complicated task since any deviation from pure spherical shape of the particles would affect their rolling behavior much more than their sliding behavior. To understand the reason, let us consider a small particle rolling on a big one. If the small particle is non spherical, the rolling motion would cause the distance between the two centers of mass to change with time, a kind of motion that requires more energy than the rolling of spherical particles that does not involve any changes in distance.
The total torque on particle $i$ due to the contact with particle $j$ is then:
$\vec{T}_{ij}=\frac{R_{i}}{R_{i}+R_{j}}\vec{r}_{ij}\times(\vec{F}_{i,friction}+\vec{F}_{i,rollingFriction})$ (8)
and the equation for the evolution of the angular velocity is then:
$\frac{2}{5}m_{i}R_{i}^{2}\frac{d\omega_{i}}{dt}=\sum_{i\neq j}\frac{R_{i}}{R_{i}+R_{j}}\vec{r}_{ij}\times(\vec{F}_{i,friction}+\vec{F}_{i,rollingFriction})$ (9)
The friction force is equal and opposite for particle $j$, so the necessary information to move both interacting particles is known from a single calculation. By dividing the torque by the respective moment of inertia, the angular acceleration applied due to surface friction is found.
2.5 Code Stability
Figure 3: Comparison between our results and analogous simulations carried out by Leinhardt & Stewart (2012). Each panel shows results for collision outcome between a projectile and a target with mass ratio $m_{p}/m_{t}$ for a specific collision angle $\theta$. Larger $\theta$ values are for grazing collisions, while $\theta=0$ represents an head-on collision. The x-axis shows the ratio of the total kinetic over total potential energies in the center of mass reference frame.
In order for the code to remain stable and conserve energy, a sufficiently large elasticity constant and a sufficiently small time step need to be used. The requirements are that (i) the maximum compression $\Delta$ attainable in the simulation needs to be a small fraction $\epsilon_{\Delta}$ of the radius of the involved particles, and that (ii) at each time step $\delta{t}$ the movement of the fastest particle is only a small fraction $\epsilon_{t}$ of the maximum compression $\Delta$. For two particles $i$ and $j$ colliding with velocity $v_{i}$ and $v_{j}$ in their centers of mass, the compression can be written as:
$\Delta_{ij}=\frac{1}{2}\frac{m_{i}v_{i}^{2}}{k}\left(1+\frac{m_{i}}{m_{j}}\right)$ (10)
In order to ensure stability (i.e., that particles cannot penetrate each-other) we require:
$\Delta_{ij}\ll\frac{r_{i}+r_{j}}{2}$ (11)
Rearranging the various terms and using $m_{i}=4\pi/3\,\rho r_{i}^{3}$ we obtain:
$k\gg 2\frac{\frac{4\pi}{3}\rho r_{i}^{3}v_{i}^{2}\left(1+\frac{m_{i}}{m_{j}}\right)}{r_{i}^{2}\left(1+\frac{r_{j}}{r_{i}}\right)^{2}}$ (12)
If we assume that $i$ is the lighter particle, then $\left(1+\frac{m_{i}}{m_{j}}\right)\leq 2$ and $\left(1+\frac{r_{j}}{r_{i}}\right)\geq 2$ and we conclude that the condition on the elasticity constant is:
$k\gg\frac{4\pi}{3}\rho r_{\max}v_{\max}^{2}$ (13)
where $r_{\max}$ is the radius of the biggest particle in the simulation and $v_{\max}$ is the speed of the fastest particle in the simulation. In most cases, the above condition yields a minimum constant for physical reliability that is much smaller than the actual Young modulus of rocks, which is of the order of $10^{11}$ barye.
The condition on the time step is inextricably linked to the chosen elasticity constant, since larger constant yields more stiff particles and therefore smaller compression, which require a smaller time step to handle. The condition we enforce is
$v_{\max}\delta{t}\ll\Delta_{ij}$ (14)
which, after analogous calculations, yields
$\delta{t}\ll\sqrt{\frac{m_{\min}}{k}}$ (15)
where $m_{\min}$ is the mass of the smallest particle in the simulation.
It is to be noted that in simple binary collisions the actual value of the elasticity constant adopted affects predominantly the duration of a particle-particle bounce. It has, instead, a negligible impact on the forces involved and on the collision outcome. It is therefore not necessary to use the actual elasticity of the material in order to obtain reliable results. For that reason, it is more productive to select a value that satisfies Eq. 13 and then use Eq. 15 to derive the required time step.
2.6 Rubble-pile Formation
For the purposes of this paper, a rubble-pile is defined as a cluster of particles in a stable structure under self-gravity. To initialize such a cluster, particles are uniformly seeded in a spherical volume. A collision/overlap check is run across all particle pairs and if a collision/overlap is detected, one of the two particles is moved randomly to another location within the spherical domain limit. This process is repeated with a tolerance of 200 tries for all particles. If any collisions/overlaps are detected after 200 tries, the sphere radius is increased by the diameter of the largest particle. All particle positions are then randomly re-seeded in the new volume and the process is re-started. This eventually leads to a reasonably well packed cluster of particles (see the left panel of Figure 1). The loosely packed particles are then allowed to collapse under their own gravity into a stable structure. A rubble-pile formation is considered complete when the kinetic energy of the system has reduced to a negligible value compared to the collapsing phase. An example of a fully formed rubble pile is shown in the right panel of Figure 1.
2.7 Cluster-finding algorithm
A post-processing code was developed to compute the outcome of the collision in term of clusters of particles with long-term stability. This action is not trivial, since the simulation cannot be carried on within a reasonable time until all particles are either unbound or part of a stable cluster. The cluster finding algorithm works iteratively in three steps. A first step runs an adaptive Gaussian smoothing filter with a local standard deviation equal to ten times each particle radius and identifies the peak of the filtered mass distribution as the center of a potential stable cluster. The second step cycles through each particle that is not already part of the potential cluster deciding whether the particle is a member of the cluster by analyzing whether or not it is in contact with a particle that belongs to the cluster. This step is repeated until no additional contacts are identified. The third step cycles again through each particle that is not already part of the potential cluster deciding whether the particle is energetically bound to the cluster. The particle velocity with respect to the center of mass of the cluster is compared to the escape velocity at the position of the particle [^3] In this step we do not consider the direction of the relative velocity nor we investigate whether the particle is in collision course with another particle. Both these can potentially alter the belonging to the cluster. . If the latter is greater, the particle is accepted into the cluster. This step is also repeated until no new particle is accepted in the cluster. At this point the whole process is repeated on all the remaining particles that are unbound from all the previously found clusters. This cycle is repeated until no new cluster with more than a user-defined minimum number of particles is found.
An example outcome is shown in Figure 2 where clusters and their member particles are color-coded. The case shown in the figure is the collision of two equal mass rubble piles with $\eta=3$. In this case six clusters with more than 50 particles are formed in the collision. The outer ones (1-green and 2-blue) are the biggest, with $M\sim 4.5\times 10^{22}$ g. The inner ones, instead, have smaller mass $M\sim 5\times 10^{21}$ g. The symmetry in the cluster properties is expected given the symmetry in the initial conditions. It has to be noted that very few of the cases in Figure 3 produce more than two cluster in the post-collision configuration. In most cases a single cluster is formed, except for the collisions between identical projectile-target, in which two similar clusters are most likely formed.
Figure 4: Cartoon showing the collision geometry between a projectile (defined ad the object with lower mass) and a target (the object with higher mass). Collisions will be parameterized on their angle $\theta$. All collisions are carried out in the center of mass reference frame in which the velocities of the two objects are anti-parallel by construction.
3 Results
We used the DECCO code to carry out three experiments. We first run a set of validation collisions, in which we aim at replicating the initial conditions from Leinhardt & Stewart (2012), repeating their collision simulations, and eventually comparing our results with theirs. Once we verified the reliability of the code, we use DECCO’s capabilities to study the effect of rubble-pile collisions on the particle size distribution and the effect of pre-collision spin of the rubble piles on the collision outcome.
3.1 Validation Runs
We test the validity of our code by carrying out a set of simulated collisions between a pair of rubble-piles of different mass and colliding at different angles. These simulations are designed to emulate the analogous ones carried out by Leinhardt & Stewart (2012). We hold the target rubble pile mass (the largest of the pair) fixed, and carry out collisions with projectiles with mass one tenth of the target, one quarter of the target, and equal to the target. For each rubble pile pair, we carry out collisions with angles $\theta=0^{\circ}$, $22^{\circ}$, and $44^{\circ}$, as defined in Figure 4. For each rubble pile pair we also carry out collisions with different relative velocity, parameterized by the ratio of the kinetic energy involved over the binding potential energy of the two rubble piles. This parameter, which we call $\eta$, is defined as:
$\eta=\frac{KE}{|U|}=\frac{\frac{1}{2}m_{p}v_{p}^{2}+\frac{1}{2}m_{t}v_{t}^{2}}{\sum_{i<j}G\frac{m_{i}m_{j}}{r_{ij}}}$ (16)
where $m_{p}$ and $m_{t}$ are the masses of the projectile and target, respectively, $v_{p}$ and $v_{t}$ are their velocities in the center of momentum frame, and $m_{i}$ and $m_{j}$ are the masses of particles $i$ and $j$. The parameter $\eta$ is allowed to vary between $0.5\leq\eta\leq 12$.
All simulations were carried out with rubble piles with a particle distribution comprising three sizes: 7.5, 15, and 22.5 km. The total mass in each particle population was equal, resulting in most of the particles being of the smaller size. All particles had density $\rho=2.7$ g/cm3 (Flynn et al., 1999). The largest (target) rubble pile had 5000 total particles, while the two smaller projectiles had 1250 and 500 total particles, respectively. These choices resulted in rubble piles with mass and radius $(m_{\rm{RP}},r_{\rm{RP}})=(6\times 10^{21},95)$, $(10^{22},129)$, and $(6\times 10^{22},203)$, in units of g and km, respectively. Impact speed ranged between 0.26 km/s for the slowest equal-mass collision to $\sim 1$ km/s for the fastest collision with the light projectile. These impact velocities are smaller than the speed of sound in dense materials, which is of the order of a few km/s.
We adopted a time step $\delta{t}=0.04$ s and elastic constant $k_{\rm{in}}=10^{18}$, with a coefficient of restitution $0.9$. The elastic constant was chosen to respect Eq. 13 for the highest energy collisions and kept constant among different simulations (even if the maximum velocity was reduced) for consistency. The time interval also comfortably satisfies the constraint set by Eq. 15. The friction coefficients that we used were $\mu_{k}=0.3$ and $\mu_{r}=10^{-3}$.
The main difference between our calculations and the ones presented in Leinhardt & Stewart (2012) is the fact that we used rubble piles with a distribution of particle sizes, while Leinhardt & Stewart (2012) use mono-dispersed particles in their experiments. We compare the mass of the largest fragment in our simulation with that reported by Leinhardt & Stewart (2012) to evaluate the reliability of the code.
Figure 3 shows the result of the validation runs. Each panel shows a rubble pile pair of a given mass ratio colliding at a given angle, and multiple relative velocities are shown. Even though subtle differences can be seen, it is clear that the two codes have similar outcomes, even in the medium-velocity regime when the fragment mass is different from either the projectile or the target. In addition, the two codes agree on the location of the sudden change in largest fragment mass that are seen in the intermediate and large angle, equal mass cases. We did not attempt to replicate their results in the high-velocity regime since our code does not include the hydrodynamic step that is used to model elastic waves and fragmentation at supersonic speed.
While other authors have carried out similar calculations with different codes, we were not able to derive all the needed parameters and outcomes from their publications, and a direct comparison was not possible.
Based on the results of this comparison, we conclude that DECCO successfully compares with previously established results and is safe to use in the subsonic regime.
3.2 Particle size distribution
Having verified that the DECCO code’s prediction compare well with the result of other codes based on analogous physics and numerical techniques, we use the specific capability of DECCO to handle particles of different sizes to investigate the role of collisions in changing the particle distributions of rubble piles. In other words, we study whether the collisions predominantly eject the largest or the smallest particles in the system. This analysis is based on the same set of simulations that were used for the validation comparison of Fig. 3. The results of the analysis are shown in Figure 5. While there is significant scattering, there are also interesting patterns. First, any effect is small, at the level of a few per cent. Repeated collisions, however, could result on a build up making the effect more noticeable. Second, most constructive collisions (blue colored symbols) and low velocity collisions tend to preferentially eject small particles, leaving a debris characterized by a bigger particle population. This is especially true for asymmetric collisions (small projectile, squares), for which all constructive collisions result in a fragment with bigger average particle size. Most destructive collisions, instead, tend to eject bigger particles preferentially, with some notable exceptions at very high velocities.
Figure 5: Effect of collisions on the average particle mass of the largest fragment. Blue symbols show cases with constructive collisions ($m_{\rm{lf}}>m_{\rm{t}}$) while orange symbols show the result of destructive collisions. Different symbols are used for systems with different projectile to target size, while the three panels show collisions with increasing collision angle.
These results can be interpreted as follows: if the energy involved in the collision is not large, it is unlikely that a single, massive particle receives enough extra energy to become unbound from the system. For this reason, constructive collisions tend to eject small particle debris and result in a stable fragment with larger average particle size. At higher energies, instead, the fact that the particles carrying the force chains are predominantly large (see Figure 6) creates a setup in which large particles can receive the bulk of the energy of the impact.
We further explore this in Figure 7. The figure is built to explore whether the high stresses are carried equally by particles of all sizes. The data shown are for an $\theta=22$ collision between a 1250 particle projectile and a 5000 particle target. Three values of the collision energy $\eta$ are shown. To create the figure, we first selected the top 1% of particles based on the strain that they are under (computed as the maximum particle overlap with all particles in contact). For each particle size $i$, the y-axis of Figure 7 shows the quantity $100\,N_{i,1\%}/N_{i}$, where $N_{i}$ is the number of particle of size $i$ in the entire simulation, and $N_{i,1\%}$ is the number of particles of size $i$ among the top $1\%$ strain particles. If all particle sizes were carrying strain equally, all y-axis values in the figures would be 1. That is not the case. As a general finding, the largest particles carry most of the strain at all times (the green line is always the largest). In addition, there is a clear evolution during and after the collision. In all cases, the pre-collision strain is emphasized with a horizontal dashed line. Let us first analyze the high-energy collision in the right panel. There is a short, few-minute spike during the collision (also shown in the right panel of Figure 6) in which large particles carry most of the strain, even more than in the pre-collision configuration. This is followed by a post-collision phase that lasts a few hours during which the rubble pile settles into a new configuration. In this phase, small grains increase their share of strain (green and orange lines are lower, blue line is higher). Eventually, the collision debris settle in a new configuration in which the largest particles still carry most of the strain but to a lesser extent with respect to the initial configuration. A similar pattern is observed for a medium energy collision in the central panel, albeit with reduced amplitude of the oscillations.
Figure 6: Volume rendering of the rubble piles before and during their collision. All particles are shown as a gray haze. Particles subject to the highest amount of stress are highlighted in dark gray. The collision is a constructive, low-velocity simulation ($\eta=3$) for a mass ratio 0.25 system with impact angle 22 degrees. The stresses are concentrated in the center of the target before the collision but the impact forces transfer the stresses to the area where the impact is happening, creating high stress chains that are carried predominantly by the bigger particles.
Figure 7: Strain analysis among the different particle sizes for a $\theta=22$, $m_{p}/m_{t}=0.25$ set of collisions of increasing energy (from left to right, $\eta=0.4$, $3.0$, and $12$). As detailed in the text, values larger than unity denote particle sizes that carry a larger than equal amount of the stress. The figure shows that large particles always carry the largest share of the stress to counter self-gravity in the clusters, and that this imbalance is even more pronounced during high-velocity impacts.
3.3 Effect of rotation
A new dedicated set of simulations was carried out to investigate the role of rotation on the collision outcome. Introducing rotations opens up a large parameter space in terms of the angular velocity and orientation of the angular momentum vectors for both the projectile and target clusters. To limit the number of simulations, we investigate only the case of head-on ($\theta$=0) collisions between same-size clusters (5000 particles each). Four different values of the collision energy were investigated ($\eta=3$, 6, 9 and 12). The angular velocity of the clusters was always set up as half of the breakup speed, and the angular momentum vector was either parallel or anti-parallel to one of the three Cartesian axes. The collision velocities are always along the x-axis. The rotational energy of the rubble piles was not included in the calculation of $\eta$, whose value is based on the center of mass kinetic energy of non-rotating analogues.
Figure 8 shows the outcome of the simulations in terms of the mass of the largest fragment. Simulations with similar angular velocity configuration were consolidated in a single data point. For example, a simulation in which the projectile’s angular momentum is along the x-axis direction and the target’s momentum is along the z-axis direction would fall in the ”perpendicular spin” case, as well as projectile’s angular momentum along the y-axis and target along the x-axis and so on. In the cases in which more than one simulation is consolidated in a single data point, the error bar shows the dispersion of the various realizations. A particularly relevant case is the one of parallel spins, since planetesimals embedded in an accretion disk are expected to be spun up by the disk differential rotation with parallel angular velocities.
Figure 8: Effect of pre-collision spinning on collision outcome for head-on collisions ($\theta=0$).
Our limited investigation in the effect of spin on the collision outcome reveals that anti-parallel spin have a very small effect, while both parallel and perpendicular spin make the collision more disruptive (the largest remmnants’ masses being smaller). This could be due to the fact that the surfaces of rubble piles colliding with either perpendicular or parallel spins have a strong sheer velocity, while rubble piles with anti-parallel spins would roll on each other. The latter is, however, true only because the two rubble piles have the same size and same angular velocity, in our case. A more thorough investigation of the effect of spin is needed to generalize these results.
4 Summary, discussion, and conclusions
We have presented a novel implementation of a discrete-element, soft sphere code for computing the dynamics of self-gravitating rubble piles. The code allows for particles of different sizes, includes inelastic normal forces between particles, gravity, kinetic friction, and rolling friction. Inelastic collisions are implemented in such a way to ensure constant coefficient of restitution, irrespective of particle size. Contact forces are not included at this stage, but are expected to be irrelevant at the particle dimensions investigated here. As the code is expanded to allow for simulation of small particles (e.g., dust grains) Van der Waals forces will be included (Kolanz et al., in preparation).
Validation runs for the code have been successfully carried out to test its reliability by reproducing previous results from analogous, well-established codes (Figure 3; Leinhardt & Stewart 2012). Given the proven validity of the code we have carried out two experiments. In the first one we have studied the particle size composition of the largest remnant from rubble pile collisions, looking for patterns of change. We found that low-velocity collisions tend to eject predominantly small particles, resulting in remnants with larger average particle size. We note that our input rubble piles do not have a segregated particle distribution with smaller particles on the surface. For this reason, more realistic simulations may find an even larger effect. For large velocity impacts that leave behind small fragments, instead, we find that the average particle size is smaller. We ascribe this behavior to the fact that the larger particles appear to be subject to the largest amount of stress, creating the force chains that propagate the force of the impact through the rubble piles (see Figures 6 and 7). Studying the strain distribution among particles of different sizes also led us to note that the post-collision configuration of the clusters is not congruent with their initial condition. The smallest particles carry an increased fraction of the strain in the post-collision configuration. This finding emphasizes the fact that a better approach to building initial cluster configurations is by hyerarchically colliding smaller clusters rather than by a monolithic collapse as described in Sect. 2.6. We plan to investigate the effect of repeated collision in a future publication.
In a second experiment we simulated collisions between fast-spinning rubble piles. We investigated the case of two identical rubble piles colliding head on ($\theta$=0). We found that collisions between rotating rubble piles lead to smaller remnants for most rotation directions. An exception was the case in which the rotation produces the effect of the two rubble piles rolling on each other. This reduced remnant size effect would be relevant in particular for collisions within accretion disks, in which the angular velocity vectors of the projectile and target objects are expected to be aligned.
While DECCO is not the only code able to carry out the simulations described here, it is important to assess the robustness of numerical results by using different codes with their own intrinsic strength and weaknesses. Future plans in term of both science and development involve adding contact forces to allow for simulations of small particles, parallelizing the code to allow for a larger number of particles, and making the code GPU portable to increase efficiency.
Acknowledgements.
This project was supported by NASA APRA award 80NSSC19K0330.